# Kinetic theory of point vortices at order ${1/N}$ and ${1/N^{2}}$

### v1

@@ III.4 Relaxation rate
We are now set to compare the prediction from Eq. (10) with measurements in $N$-body simulations. This is presented in Fig. 5.
@@
Figure 5: Initial relaxation rate in action space, $\mathcal{R}_{1}$, for profile (1). For the simulations, the uncertainty is illustrated with the 16% and 84% contours obtained by bootstrap over the available realisations (see Appendix G). For the prediction, the full line corresponds to the ${1/N}$ Landau equation (Eq. 10), while the dashed line is from the “regularised” ${1/N}$ Landau equation (Eq. 25), with the regularised time, ${T_{\mathrm{reg}}/T_{\mathrm{dyn}}\!=\!307}$, adjusted to the slope around the sign switch of $\mathcal{R}_{1}$. Close to the extremum of the frequency profile, broadening effects play a crucial role.
@@
A couple of remarks can be made on Fig. 5. For now, let us focus on the ${1/N}$ kinetic prediction given by Eq. (10). First, for ${J/J_{0}\!\lesssim\!0.9}$, it predicts ${\mathcal{R}_{1}\!=\!0}$. Glancing back at Fig. 2, this corresponds to a range of actions, $J$, for which the resonance condition, ${\Omega(J_{1})\!=\!\Omega(J)}$ (Eq. 14), imposes ${J_{1}\!=\!J}$. As such, only local resonances are permitted by the mean frequency profile, hence leading to the local vanishing of Eq. (14) because of the antisymmetry of the collision kernel. Therefore, at order ${1/N}$, this region does not relax towards the Boltzmann distribution, because of a local kinetic blocking. As such, to relax towards the statistical equilibrium state, $1/N^{2}$ effects must be taken into account, as studied in the next section for a purely monotonic frequency profile. Then, for ${J/J_{0}\!\simeq\!1.03}$, we find that the kinetic prediction for $\mathcal{R}_{1}$ abruptly changes sign. This is due to the extremum of the frequency profile that appears at this location, as visible in Fig. 2. Although the kinetic prediction is discontinuous near this extremum, it suffers from no divergences.
@@
We can now compare this prediction with the measurement from the $N$-body simulations. Reassuringly, we recover that for ${J/J_{0}\!\lesssim\!0.9}$, relaxation is greatly delayed, a local kinetic blocking. Second, as one considers larger action, say ${0.9\!\lesssim\!J/J_{0}\!\lesssim\!1.0}$, we find a good agreement between the numerical simulations and the kinetic prediction from the ${1/N}$ Landau equation. However, as ${J}$ nears ${1.03\,J_{0}}$, i.e. as $J$ nears the extremum of the frequency profile, we find that the kinetic prediction overestimates the numerical measurements. In particular, the simulations do not exhibit the discontinuity predicted by the Landau equation, and present a rather smooth change of sign. This mismatch is to be attributed to the limitations of some of the hypotheses made in the derivation of the ${1/N}$ Landau equation. In Fig. 5, we present a prediction with a “regularised” Landau equation, that offers a much more satisfying agreement with the numerical simulation. This is the topic of the next section.
@@ III.5 Resonance broadening
The ${1/N}$ Landau equation (Eq. 14) involves the sharp resonance condition, ${\delta_{\mathrm{D}}[\Omega(J)\!-\!\Omega(J_{1})]}$. Mathematically speaking, this sharp resonant Dirac comes from the fact that the derivation of the Landau equation requires the propagation of the dynamics of the DF’s fluctuations, at linear order, up to infinite dynamical times (see, e.g., Chavanis, 2012c). Yet, as a result of non-linearities and diffusion, such an assumption is only partially true. This effect can be responsible in particular for “resonance broadening” (see, e.g. Weinstock, 1969), as we will now briefly explore, in a heuristic fashion.
[…]
@@
with ${\omega_{\mathrm{R}}\!=\!\Omega(J)\!-\!\Omega(J_{1})}$ the resonance frequency. Therefore, in Eq. (21), we recover, heuristically, the sharp resonance condition from Eq. (14). To sum up, the ${1/N}$ Landau equation involves a Dirac delta as a result of two assumptions (i) the DF fluctuations are propagated linearly, i.e. along the unperturbed orbits; (ii) and this holds for infinitely long dynamical times.
Yet, these two assumptions cannot hold for so long, since we neglected the contributions from the nonlinear term, ${[\delta F,\delta H]}$, in Eq. (19). Heuristically, let us therefore assume that Eq. (20) must be weighted down by an exponential damping factor on some “regularisation” timescale, $T_{\mathrm{reg}}$. The dynamics of the DF fluctuations would become
@@
Naturally, in the limit ${T_{\mathrm{reg}}\!\to\!+\infty}$, one has ${\delta_{T_{\mathrm{reg}}}(\omega)\!\to\!\delta_{\mathrm{D}}(\omega)}$, hence recovering a sharp resonance condition.
@@
[…]
Of course, we stress that Eq. (25) is only a heuristic kinetic equation, since, for example, it does not conserve the total energy, contrary to Eq. (14).
@@
In Fig. 6, we compute the relaxation rate predicted by Eq. (25), as one varies the regularisation timescale, $T_{\mathrm{reg}}$.
@@
Figure 6: Predicted relaxation rate for profile (1) given by Eq. (25), as one varies the regularisation time $T_{\mathrm{reg}}$, using the same convention as in Fig. 5. For ${T_{\mathrm{reg}}\!\to\!+\infty}$, one recovers the ${1/N}$ Landau prediction from Eq. (14).
@@
In that figure, we find that broadening effects smooth out the discontinuity predicted by the ${1/N}$ Landau equation (Eq. 14) close to the extremum of the frequency profile. Reassuringly, we also find that as ${T_{\mathrm{reg}}\!\to\!+\infty}$, the regularised prediction converges to the (sharp) Landau prediction.

### v2

@@ III.4 Relaxation and diffusion
We are now set to compare the prediction from Eq. (10) with measurements in $N$-body simulations. This is presented in Fig. 5, where we investigate both the relaxation rate and the diffusion coefficient.
@@
Figure 5: Initial relaxation rate in action space, $\mathcal{R}_{1}(J)$} (Eq. 20a, top panel), and diffusion coefficient, $\mathcal{D}(J)$} (Eq. 20b, bottom panel), for profile (1). For the simulations, the uncertainty is illustrated with the 16% and 84% contours obtained by bootstrap over the available realisations. We refer to Appendix G for details on the numerical setup. For the prediction, the full line corresponds to the 1/N{1/N} Landau equation (Eqs. 10 and 17), while the dashed line is from the “regularised” 1/N{1/N} Landau equation (Eqs. 32 and 33), with the regularised time, $T_{\mathrm{reg}}/T_{\mathrm{dyn}}\!=\!307$}, adjusted to the slope around the sign switch of $\mathcal{R}_{1}(J)$}. Close to the extremum of the frequency profile, broadening effects play a crucial role.
A couple of remarks can be made on Fig. 5.
III.4.1 Analytical predictions
@@
For now, let us focus solely on the kinetic predictions, for the Landau flux (Eq. 10) in the top panel and the diffusion coefficient (Eq. 17) in the bottom panel. First, for $J/J_{0}\!\lesssim\!0.9$}, it predicts $\mathcal{R}_{1}(J)\!=\!0$}, while, for the same range of action, one finds $\mathcal{D}(J)\!>\!0$}. Glancing back at Fig. 2, this corresponds to a range of actions, $J$, for which the resonance condition, $\Omega(J_{1})\!=\!\Omega(J)$} (Eq. 14), imposes $J_{1}\!=\!J$}. As such, only local resonances are permitted by the mean frequency profile. These local resonances lead to the local vanishing of Eq. (14) because of the antisymmetry of the collision kernel. On the contrary, such local resonances always contribute to the diffusion coefficient. As a result, the diffusion coefficient is non-zero over the whole action domain considered. At order 1/N{1/N}, because its overall relaxation rate vanishes, the action region, $J/J_{0}\!\lesssim\!0.9$} does not relax towards the Boltzmann distribution. This is a local kinetic blocking. As such, to relax towards the statistical equilibrium state, $1/N^{2}$ effects must be taken into account. This will be studied in Section IV for a purely monotonic frequency profile.
Then, in Fig. 5, for $J\!\to\!J_{\star}\!\simeq\!1.03\,J_{0}$}, we find that the kinetic prediction for $\mathcal{R}_{1}(J)$} abruptly changes sign, while the prediction for $\mathcal{D}(J)$} diverges locally. This is due to the extremum of the frequency profile in $J\!=\!J_{\star}$}, as visible in Fig. 2. Let us now detail the origin of such a behaviour.
For $J\!\to\!J_{\star}$}, we may, up to a constant, approximate the frequency profile with the quadratic dependence $\Omega(J)\!\simeq\!(J\!-\!J_{\star})^{2}$}, see Fig. 2. As a result, when evaluated in $J\!=\!J_{\star}\!+\!\varepsilon$}, Eqs. (14) and (17) predict for the relaxation rate and diffusion coefficient a behaviour of the form
$\mathcal{R}_{1}(J_{\star}\!+\!\varepsilon){}$ $\propto\!\!\int\!\!\mathrm{d}J_{1}\,\delta_{\mathrm{D}}\big[\varepsilon^{2}\!-\!\big(J_{1}\!-\!J_{\star}\big)^{2}\big]$ (22a)
$\times{}$ $\big[F^{\prime}(J_{\star}\!+\!\varepsilon)F(J_{1})\!-\!F(J_{\star}\!+\!\varepsilon)F^{\prime}(J_{1})\big],$
$\mathcal{D}(J_{\star}\!+\!\varepsilon){}$ $\propto\!\!\int\!\!\mathrm{d}J_{1}\,\delta_{\mathrm{D}}\big[\varepsilon^{2}\!-\!\big(J_{1}\!-\!J_{\star}\big)^{2}\big]\,F(J_{1}).$ (22b)
where we used the notation $F^{\prime}(J)\!=\!\partial F/\partial J$}, and dropped the coupling coefficient, $|\Lambda_{\mathrm{tot}}|^{2}$}, taking it to be constant.
To perform the integral over action space, we use the generic relation
$\delta_{\mathrm{D}}[f(J)]=\sum_{J_{\mathrm{r}}}\delta_{\mathrm{D}}[J\!-\!J_{\mathrm{r}}]/|\Omega^{\prime}(J_{\mathrm{r}})|,$ (23)
where the sum runs over the resonant actions, $J_{\mathrm{r}}$, such that $f(J_{\mathrm{r}})\!=\!0$}. In the present case, the resonance condition from Eq. (22) imposes $J_{1}\!=\!J_{\star}\!\pm\!\varepsilon$}. As previously argued, we find that the contribution from the local resonance $J_{1}\!=\!J_{\star}\!+\!\varepsilon\!=\!J$} does not contribute to the flux, $\mathcal{R}_{1}(J)$}, while it contributes to the diffusion coefficient, $\mathcal{D}(J)$}. The non-local resonance $J_{1}\!=\!J_{\star}\!-\!\varepsilon$} contributes to both quantities. Equation (22) becomes
$\mathcal{R}_{1}(J_{\star}\!+\!\varepsilon){}$ $\propto\frac{1}{|\varepsilon|}\,\big[F^{\prime}(J_{\star}\!+\!\varepsilon)F(J_{\star}\!-\!\varepsilon)\!-\!F(J_{\star}\!+\!\varepsilon)F^{\prime}(J_{\star}\!-\!\varepsilon)\big],$ (24a)
$\mathcal{D}(J_{\star}\!+\!\varepsilon){}$ $\propto\frac{1}{|\varepsilon|}\,\big[F(J_{\star}\!+\!\varepsilon)+F(J_{\star}\!-\!\varepsilon)\big].$ (24b)
In the limit $\varepsilon\!\to\!0$}, i.e. as one nears the extremum of the frequency profile, one finds the asymptotic behaviours
$\mathcal{R}_{1}(J_{\star}\!+\!\varepsilon){}$ $\simeq\mathrm{Sign}[\varepsilon]$ $\!\!\!\!\!\!\!\!\text{for}\quad\varepsilon\to 0,$ (25a)
$\mathcal{D}(J_{\star}\!+\!\varepsilon){}$ $\simeq\frac{1}{|\varepsilon|}$ $\!\!\!\!\!\!\!\!\text{for}\quad\varepsilon\to 0.$ (25b)
As such, as visible in Fig. (5), the 1/N{1/N} Landau equation predicts that (i) the flux in action space is discontinuous near the extremum of the frequency profile, but suffers from no divergence there; (ii) the diffusion coefficient in action space suffers from a genuine local divergence.
III.4.2 Numerical measurements
@@
We can now compare the predictions of the flux and diffusion coefficients with measurements in $N$-body simulations. First, for the flux (top panel of Fig. 5), we recover that for $J/J_{0}\!\lesssim\!0.9$}, relaxation is greatly delayed. This is the imprint of the local kinetic blocking. Second, as one considers larger action, say $0.9\!\lesssim\!J/J_{0}\!\lesssim\!1.0$}, we find a good agreement between the numerical simulations and the kinetic prediction, for both the flux and the diffusion coefficient. However, as J{J} nears $J_{\star}$, the location of the extremum of the frequency profile, we find that the kinetic prediction overestimates the numerical measurements. Indeed, the simulations do not exhibit the discontinuity of the flux nor the divergence of the diffusion coefficient, predicted by the 1/N{1/N} Landau equation. In particular, the numerically-measured flux, $\mathcal{R}_{1}(J)$}, exhibits a rather smooth change of sign, while the numerically-measured diffusion coefficient, $\mathcal{D}(J)$}, presents a two-hump shape that only reaches some finite amplitude. Finally, further out in action space, i.e. for $J/J_{0}\!\gtrsim\!1.1$}, we find that the kinetic prediction for the flux matches with the $N$-body simulations, while for the diffusion coefficient, some mismatch between the two remains. We note that this domain corresponds to actions, $J$, whose orbital frequency, $\Omega(J)$}, is still close to the extremum one, $\Omega(J_{\star})$}.
The mismatch observed in Fig. 5 is to be attributed to the limitations of some of the hypotheses made in the derivation of the 1/N{1/N} Landau equation. In Fig. 5, we present a prediction with a “regularised” Landau equation. It offers a much more satisfactory agreement with the numerical simulations. For the diffusion coefficient, the match is improved, though far from ideal. Nonetheless, this heuristic theory still manages to reproduce the smooth change of sign of the flux and the two-hump structure of the diffusion coefficient. The exploration of such “regularised” kinetic predictions is the topic of Section III.5.
III.4.3 Comparison with other 1D{1D} systems
Naturally, one could wonder which specific features of the current setup make it more akin to such discontinuities and the need for regularisation. Indeed, it was shown for 1D{1D} gravity (Roule et al., 2022) and the Hamiltonian Mean Field model (HMF) model (Benetti and Marcos, 2017) that the 1/N{1/N} Landau equation can provide an excellent description of 1D{1D} relaxation processes. We briefly review here some of the key differences between these systems.
In the case of 1D{1D} gravity, there is no divergence on small-scale of the kinetic prediction (see, e.g., eq. (A7) in Roule et al., 2022). In addition, the frequency profile of a typical stable quasistationary state exhibits no extremum (see fig. 1 in Roule et al., 2022). Physically, this means that the contribution from phase mixing never vanishes, so that there is always a non-zero shear in action space. For these two reasons, the 1/N{1/N} Landau equation produces no local discontinuities or divergences in action space. For that system, one finds then an excellent match between the 1/N{1/N} kinetic prediction and numerical measurements, both for the diffusion coefficient (see fig. 2 in Roule et al., 2022) and the relaxation rate (see fig. 3 in Roule et al., 2022).
Let us now consider the case of the HMF model (Antoni and Ruffo, 1995). Since the pairwise interaction in that system involves only the $k\!=\!\pm 1$} large-scale harmonics, the 1/N{1/N} Landau equation does not suffer from any divergence on small scales. In the limit of a homogeneous quasistationary state, i.e. F=F(v){F\!=\!F(v)}, the system’s action is $J\!\sim\!v$}, and the frequency profile is simply $\Omega(J)\!\sim\!\partial_{v}[\tfrac{1}{2}v^{2}]\!\sim\!v$}, making it strictly monotonic. In the homogeneous limit, the HMF model is therefore submitted to a global kinetic blocking of its 1/N{1/N} dynamics, similarly to profile (2) here. In that case, the overall relaxation is driven only by $1/N^{2}$} effects (Rocha Filho et al., 2014). In the limit of weak collective amplification, the $1/N^{2}$} Landau equation offers a convincing match with the numerical measurements (see fig. 2 in Fouvry et al., 2019).
Finally, let us consider the case of the HMF model when in an inhomogeneous, i.e. magnetised, quasistationary state. Here again, there is no divergence on small scales of the 1/N{1/N} kinetic theory. The typical frequency profile of such a system exhibits two striking features (see, e.g., fig. 3 in Benetti and Marcos, 2017): (i) For $J\!\to\!0$}, i.e. in the centre of the trapped region, one has $\Omega^{\prime}(J)\!\to\!0$}; (ii) For $J\!\to\!J_{\mathrm{sep}}$}, i.e. at the system’s separatrix, one has $|\Omega^{\prime}(J)|\!\to\!+\infty$}.
Naively, one could fear that the vanishing of the frequency gradient in the system’s centre would lead to the same asymptotic behaviour as in Eq. (25). In practice, for the inhomogeneous HMF model, one finds that the pairwise coupling coefficient satisfies $\lim_{J\to 0}U_{k}(J,J)\!=\!0$} (see eqs. (45) and (46) in Benetti and Marcos, 2017). Physically, this makes sense since orbits in the system’s centre correspond to particles that are exactly at rest. As a result, the Fourier transform in angle of the interaction potential between two such orbits therefore vanishes (see eq. (44) in Benetti and Marcos, 2017). This differs greatly from the present case of point vortices where, at the extremum of the frequency profile, one finds that $\lim_{J\to J_{\star}}U_{k}(J,J)$} is non-zero. Because of the vanishing of the pairwise coupling coefficient in the centre of the inhomogeneous HMF model, one finds that both the relaxation rate and the diffusion coefficient vanish there. Therefore, the associated 1/N{1/N} kinetic prediction exhibits no discontinuities nor divergences, and matches with numerical measurements (see, e.g., fig. 6 in Benetti and Marcos, 2017).
At the location of the HMF separatrix, we note that the gradient of the frequency profile tends to infinity. Physically, this means that close the separatrix, pairwise resonances are swept extremely fast, rather than extremely slow as is the case close to a vanishing frequency gradient. As such, the type of discontinuities and divergences emphasised in Eq. (25) are not to be expected close to the HMF separatrix. Nonetheless, we note that this rapidly varying frequency profile leads to localised, but finite, “spikes” in the prediction for the Landau diffusion coefficient. These are clearly visible in fig. 2 of Fouvry and Bar-Or (2018) and are in agreement there with $N$-body measurements. Finally, we point out that the non-monotonicity of the frequency profile of the inhomogeneous HMF model is essential to prevent this system from suffering from a global kinetic blocking. Indeed, it is the non-local resonant couplings between particles located on each side of the separatrix that are solely responsible for driving the overall 1/N{1/N} relaxation.
@@ III.5 Resonance broadening
The 1/N{1/N} Landau equation (Eq. 14) involves the sharp resonance condition, $\delta_{\mathrm{D}}[\Omega(J)\!-\!\Omega(J_{1})]$}. Mathematically speaking, this sharp resonant Dirac comes from the fact that the derivation of the Landau equation requires the propagation of the dynamics of the DF’s fluctuations, at linear order, up to infinite dynamical times (see, e.g., Chavanis, 2012c). Yet, as a result of nonlinearities and diffusion, such an assumption is only partially true. This effect can be responsible in particular for “resonance broadening” (see, e.g. Weinstock, 1969), as we will now explore, in a heuristic fashion.
III.5.1 Heuristic regularisation of the Landau equation
Typically, the derivation of the Landau equation starts with the introduction of the system’s empirical DF, $F_{\mathrm{d}}\!=\!\sum_{i}\gamma_{i}\delta_{\mathrm{D}}[\mathbf{r}\!-\!\mathbf{r}_{i}(t)]$}. This DF evolves according to the Klimontovich equation, $\partial_{t}F_{\mathrm{d}}\!+\![F_{\mathrm{d}},H_{\mathrm{d}}]\!=\!0$}, with $H_{\mathrm{d}}\!=\!H_{\mathrm{d}}[F_{\mathrm{d}}]$} the system’s (specific) empirical Hamiltonian, and $\cdot$$\cdot$]} the Poisson bracket (Chavanis, 2012c). The empirical DF is then expanded into $F_{\mathrm{d}}\!=\!F\!+\!\delta F$}, with $F$ the (smooth) ensemble-averaged DF from Section II.3, and $\delta F$} its fluctuations. The crucial step of the quasilinear assumption is to assume that the dynamics of $\delta F$} can be expanded at linear order, to become
@@ III.5.1 Heuristic regularisation of the Landau equation
with $\omega_{\mathrm{R}}\!=\!\Omega(J)\!-\!\Omega(J_{1})$} the resonance frequency. Therefore, in Eq. (28), we recover, heuristically, the sharp resonance condition from Eq. (14). To sum up, the 1/N{1/N} Landau equation involves a Dirac delta as a result of two assumptions (i) the DF fluctuations are propagated linearly, i.e. along the unperturbed orbits; (ii) and this holds for infinitely long dynamical times.
Yet, these two assumptions cannot hold for so long, since we neglected the contributions from the nonlinear term, $\delta F,\delta H]$}, in Eq. (26). Heuristically, let us therefore assume that Eq. (27) must be weighted down by an exponential damping factor on some “regularisation” timescale, $T_{\mathrm{reg}}$. The dynamics of the DF fluctuations would become
@@
Naturally, in the limit $T_{\mathrm{reg}}\!\to\!+\infty$}, one has $\delta_{T_{\mathrm{reg}}}(\omega)\!\to\!\delta_{\mathrm{D}}(\omega)$}, hence recovering a sharp resonance condition. We refer to Appendix I.1 for an alternative approach to justify the appearance of the broadened Dirac from Eq. (31) by taking into account the stochasticity of the particles trajectories.
@@
$\times{}$ $\,\big|\Lambda_{\mathrm{tot}}(J,J_{1})\big|^{2}\,\bigg(F(J_{1})\,\frac{\partial F}{\partial J}\!-\!F(J)\,\frac{\partial F}{\partial J_{1}}\bigg)\bigg].$
Similarly, following the same line of thought, the diffusion coefficient from Eq. (17) can be regularised into
$D_{2}(J)=2\gamma\!\!\int\!\!\mathrm{d}J_{1}\,\delta_{T_{\mathrm{reg}}}\big[\Omega(J)\!-\!\Omega(J_{1})\big]\big|\Lambda_{\mathrm{tot}}(J,J_{1})\big|^{2}\,F(J_{1}).$ (33)
As a side note, we point out that the broadened Dirac in Eq. (31) is always positive. This ensures that the regularised diffusion coefficient from Eq. (33) is always positive. This is an important requirement to comply with the physical definition of the diffusion coefficient from Eq. (18). Of course, we stress that Eqs. (32) and (33) are only heuristic expressions. Surely, they would deserve further theoretical developments. For example, Eq. (32) does not conserve the total energy, contrary to Eq. (14). Some additional properties of the regularised Landau equation (Eq. 32) are also discussed in Appendix H, in particular how it regularises both the generic divergence of the diffusion coefficient on small scales (Appendix H.1), and the divergence of the diffusion coefficient at the extremum of the frequency profile (Appendix H.2).
@@
In Fig. 6, we compute the regularised relaxation rate and diffusion coefficients predicted by Eqs. (32) and (33), as one varies the regularisation timescale, $T_{\mathrm{reg}}$.
@@
Figure 6: Predicted initial regularised relaxation rate (Eq. 32, top panel) and diffusion coefficient (Eq. 33, bottom panel), for profile (1), as one varies the regularisation time $T_{\mathrm{reg}}$, using the same convention as in Fig. 5. For $T_{\mathrm{reg}}\!\to\!+\infty$}, one recovers the 1/N{1/N} Landau prediction from Eqs. (14) and (17).
@@
In that figure, we find that broadening effects smooth out the discontinuity and local divergence predicted by the 1/N{1/N} Landau equation (Eq. 14) close to the extremum of the frequency profile. Reassuringly, we also find that as $T_{\mathrm{reg}}\!\to\!+\infty$}, the regularised prediction converges to the (sharp) Landau prediction. As such, varying $T_{\mathrm{reg}}$ in Eq. (32) allows for a smooth transition between the usual 1/N{1/N} Landau equation and the nonlinear regime of resonance broadening. In Appendices H and I, we discuss this connexion further. In particular, we link our result with the ones obtained in previous literature in the regime of sheared and unsheared flows.
III.5.2 Finite-time diffusion coefficients
In order to visualise the appearance of resonance broadening in the simulations, taking inspiration from Eq. (18), let us introduce the finite-time diffusion coefficient as
$D_{2}(J,T)=\frac{\big\langle\big[\Delta J(T)\big]^{2}\big\rangle}{T}.$ (34)
In practice, the 1/N{1/N} Landau equation predicts that $\lim_{T\to+\infty}D_{2}(J,T)\!=\!D_{2}(J)$}, with the Landau diffusion coefficient given by Eq. (17). Deviations from this limit are therefore to be attributed to nonlinear contributions spoiling some of the underlying assumptions of the 1/N{1/N} Landau equation. Naturally, from $D_{2}(J,T)$}, one can follow Eq. (20b) and readily define the associated rescaled finite-time diffusion coefficient, $\mathcal{D}(J,T)$}.
In Fig. 7, we present numerical measurements of the finite-time diffusion coefficients, $\mathcal{D}(J,T)$}, for various times $T$.
Figure 7: Illustration of the finite-time diffusion coefficient, $\mathcal{D}(J,T)$} (Eq. 34), as measured in $N$-body simulations for various finite times $T$, for profile (1). The dashed line is the 1/N{1/N} Landau prediction from Eq. (17). We refer to Appendix G for details on the numerical setup. As $T$ increases, the finite-time diffusion coefficient grows, but eventually saturates to some finite amplitude, as a result of resonance broadening.
In that figure, we note that, as one increases $T$, the finite-time diffusion coefficients start by following more and more closely the 1/N{1/N} Landau diffusion coefficient. Yet, for $T/T_{\mathrm{dyn}}\!\gtrsim\!320$}, i.e. a timescale similar to the value of $T_{\mathrm{reg}}$ introduced in Fig. 5, we note that (i) the growth of $\mathcal{D}(J,T)$} halts; (ii) it starts to exhibit a two-hump structure reminiscent of the one of the regularised prediction from Fig. 6. As one pushes $T$ further, the numerically-measured finite-time diffusion coefficients even start to decay. This is because for such long time separations, the test particles have now entered their “diffusive” regime, i.e. $T\!\to\!\mathcal{D}(J,T)$} growing linearly in time, as illustrated numerically in Fig. 18 of Appendix G. In addition, we note that the mismatch between the finite-time diffusion coefficient and the 1/N{1/N} Landau prediction is the strongest for $J\!\geq\!J_{\star}$} rather than for $J\!\leq\!J_{\star}$}. This is because particles with $J\!\geq\!J_{\star}$} are the ones with an orbital frequency, $\Omega(J)$}, closer to the extremum frequency, $\Omega(J_{\star})$}, as visible in Fig. 2. Overall, we argue that two-hump structure in Fig. 7 along with the saturation to finite amplitude are numerical testimonies that resonance broadening effects close to the extremum of the frequency profile play an instrumental role in the relaxation of this system.
III.5.3 Numerical comparison